Circular No. 8498

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE SUPERNOVA IN NGC 4656 N. Elias-Rosa, S. Benetti, E. Cappellaro, and M. Dolci, Istituto Nazi-

onale di Astrofisica; and A. Pastorello, Max-Planck-Institut, report that a spectrum of the new object in NGC 4656 (cf. IAUC 8497), obtained on Mar. 23.9 UT by W. Boschin and N. Pinilla Alonso with the TNG telescope (+ Dolores; range 350–800 nm, resolution 1.2 nm), shows a relatively blue continuum dominated by a narrow (730 km/s) Hα emission, while the other Balmer lines are not visible. The broader component of the Hα line often seen in type-IIn supernovae is not apparent. The spectrum also shows narrow (1.5-nm) absorptions due to H and K of Ca II. This object may be a superoutburst of a luminous blue variable rather than a true supernova, similar to SN 1997bs (IAUC 6627; Van Dyk et al. 2000, PASP 112, 1532) and 1999bw (IAUC 7152).

V382 NORMAE L. A. G. Monard revises his position for V382 Nor (cf. IAUC 8497) to

α = 16^{h}19^{m}44^{s}.74, δ = −51^{o}34′53′′.1 (equinox 2000.0). Visual magnitude estimates: Mar. 20.195 UT, 10.1 (R. Y. Shida, São Paulo, Brazil); 20.880, 10.1 (A. Pearce, Nedlands, W. Australia); 22.615, 10.2 (Pearce).

CATACLYSMIC VARIABLE IN PYXIS Corrigendum. On IAUC 8495, line 6, for H. Haseda, read K. Haseda, Visual magnitude estimate from A. Pearce, Nedlands, W. Australia:

Mar. 20.546 UT, 13.3.

COMETS C/2005 C1, C/2005 C2, C/2005 C3, AND C/2005 C4 (SOHO)

The initial observations are given below for additional comets found on SOHO website images (cf. IAUC 8494). C/2005 C2 and C/2005 C3 are Kreutz sungrazers, while C/2005 C1 belongs to the Meyer group; C/2005 C4 seems not to be a group member.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 C1 Feb. 2.131 21^{h}07.^{m}9 −15^{o}30′ C2  HS 2005-F32
C/2005 C2 2.138 21 18.7 −18 13 C3 HS 2005-F32
C/2005 C3  7.179 21 52.1 −17 39 C3/2 XL 2005-F32
C/2005 C4  12.704 21 48.8 −14 44 C2  RK 2005-F32

2005 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green